OpenStudy (anonymous):

Solve: P=2(l+w) (solve for l) I got P-w=l

OpenStudy (anonymous):

First Divide By 2 p/2=l+w Second subtract w p/2-w=l

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

Yw :)